### v1

# Bone: Block Affine Transformation as Parameter Efficient Fine-tuning Methods for Large Language Models

Abstract
As Large Language Models (LLMs) continue to grow in size, their computational and memory requirements increase correspondingly. Consequently, the exploration of cost-effective and efficient fine-tuning methods has become increasingly important. Low-Rank Adaptation (LoRA) has achieved remarkable training results by freezing the original weights and training only low-rank matrices, establishing itself as the predominant fine-tuning method for LLMs. In pursuit of performance closer to full-parameter training, a series of LoRA variants have emerged, such as LoRA+, PISSA, Olora, and LoRA-GA. However, these methods also make the fine-tuning initialization process more complex, and it remains challenging to surpass the performance ceiling of full fine-tuning. To address these issues, this paper introduces an innovative method called Bone (Block Affine), which not only reduces memory overhead but also emphasizes the internal connections between weights, leading to faster convergence and better data fitting. Experimental comparisons across two different LLM architectures (LLaMA2, RWKV6) and various parameter scales demonstrate that the Bone structure can achieve rapid convergence and superior data fitting without the need for complex initialization. For example, when fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8k and math benchmarks, Bone achieved fine-tuning scores of 49.36 and 8.8, respectively, outperforming PISSA by 5.84% and 1.96%.
1 Introduction
Large models have been integrated into various industries, revolutionizing many traditional technologies Radford et al. (2019); Raffel et al. (2020). However, general-purpose large models often struggle to meet the needs of all downstream tasks, making it necessary to fine-tune base models for specific scenarios. Full-scale fine-tuning of large models is extremely costly; for example, training the LLaMA2-7B Touvron et al. (2023) model with bfloat16 Wang & Kanwar (2019) precision requires around 60GB of VRAM. In contrast, by using PEFT (Parameter-Efficient Fine-Tuning)Xu et al. (2023) techniques, fine-tuning can be done with just a single 24GB VRAM GPU. As a result, numerous PEFT techniques and quantization methods have emerged to reduce the training costs of large models. LoRA (Low-Rank Adaptation) Hu et al. (2021) has become one of the most popular PEFT methods due to its small parameter size, its ability to optimize the model’s main weights, and the fact that it doesn’t add extra inference overhead.
LoRA significantly reduces memory usage by freezing the original weights $W$ and updating two low-rank matrices $A$ and $B$. Typically, either $A$ or $B$ is initialized to zero, ensuring that the initial state of LoRA is consistent with the pre-trained model, The Figure 2(a) illustrates the structure visualization. However, extensive experiments Ding et al. (2023); Liu et al. (2024b); Biderman et al. (2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. This slow convergence is likely due to the small gradients caused by the zero initialization of either $A$ or $B$. To address this issue, researchers have proposed several LoRA variants, such as LoRA+ Hayou et al. (2024), PISSA Meng et al. (2024), and LoRA-GAWang et al. (2024). Despite their excellent performance, these LoRA variants inevitably introduce complexity into the fine-tuning process and reduce structural flexibility. For example, LoRA+ requires manual adjustment of different learning rates for $A$ and $B$; PISSA necessitates performing SVD decomposition on $W$ at the beginning of training, which can be time-consuming when the model parameters are large.
In this paper, we propose a novel PEFT method called Bone (Block Affine) to overcome the limitations of LoRA variants and surpass full fine-tuning. It is well-known that LLMs have extremely large parameters, with complex interactions among their internal weights. To calculate the information interactions within the weight matrix $W$, Bone employs an initially zeroed, trainable matrix, named bone, to perform Block Affine operations on $W$. The specific calculations of Block Affine are detailed in Section 3, and the structural visualization is shown in Figure  2. To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 Xu et al. (2023), RWKV6 Peng et al. (2024)). As shown in Figure  1, Bone achieves the fastest convergence, and the results on the test set demonstrate its superior data fitting and generalization capabilities (Table  1, Table  2).
Advantages of Bone::
1. Bone does not require the complex initialization needed by LoRA variants, yet it achieves fast convergence and better data fitting.
2. Block Affine not only efficiently promotes information exchange between weights but also enhances the weight utilization of low-rank matrices.
3. Bone provides a new direction for the development of PEFT techniques.
(a) llama2-7b
(b) RWKV6
Figure 1: The left image shows the loss curve for LLaMA2-7B fine-tuned on the MetaMathQA dataset, with the first 100 steps highlighted for closer observation. The right image displays the loss curves for RWKV6 with different parameter settings, also fine-tuned on MetaMathQA. Both models were trained with a batch size of 64 for a full epoch. Comparing the loss curves reveals that Bone demonstrates superior fitting ability across various architectures and parameter settings. Additionally, Bone exhibits a rapid decrease in loss within the first 100 steps, highlighting its effectiveness.
2 Related Works
The PEFT (Parameter-Efficient Fine-Tuning) techniques are diverse and include approaches like adapter tuningHoulsby et al. (2019), prefix tuningLiu et al. (2023), prompt tuningBrown (2020), LoRAHu et al. (2021), and layer-freezing methods such as LISA.
The adapter method does not require fine-tuning all the parameters of the pre-trained model. Instead, it introduces a small number of task-specific parameters to store knowledge related to that task, thereby reducing the computational demands of model fine-tuning. Prefix tuning is a lightweight fine-tuning method for generative tasks. It adds a continuous, task-specific vector sequence, called a prefix, to the input. Unlike prompts, prefixes are entirely composed of free parameters and do not correspond to actual tokens. Compared to traditional fine-tuning, prefix tuning only optimizes the prefix, meaning we only need to store a large Transformer model and a small set of task-specific prefixes, resulting in minimal overhead for additional tasks.
Prompt tuning defines a unique prompt for each task, appending it to the input data while freezing the pre-trained model during training. LoRA allows us to indirectly train certain dense layers in a neural network by optimizing low-rank decomposition matrices that adapt these layers, while keeping the pre-trained weights unchanged. This approach significantly addresses the inference latency introduced by adapters and the fitting challenges of prefix tuning.
In LoRA, the adapter matrices $A$ and $B$ are updated with the same learning rate, but using the same rate for both may not effectively learn the features. LoRA+ corrects this suboptimality by setting different learning rates for matrices $A$ and $B$ with a fixed ratio, improving the method’s efficiency. The DoRA Liu et al. (2024b) method combines weight decomposition to achieve learning capabilities similar to full fine-tuning without sacrificing LoRA’s inference efficiency. PiSSA optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), PiSSA initializes the dominant singular values and vectors to train these matrices, while keeping the residual matrix static during fine-tuning. OLoRA Büyükakyüz (2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $A$ and $B$ by computing the initial gradient, thereby closely approximating full fine-tuning.
3 Bone: Block Affine
In this section, we will introduce the Bone method. Unlike the LoRA series of architectures, Bone employs a single learnable matrix, initialized to zero and named ”bone.” This matrix interacts with the model’s matrix $W$ through Block Affine operations to represent $\Delta W$. This approach ensures that the initial state remains unchanged while also facilitating the integration of information between the weights. Assuming the matrix $W$ has dimensions $(n,n)$, ideally, we train with a full-rank matrix $bone$ $(n,n)$. The specific calculation for Block Affine is as follows:
$Y=X(W+\Delta W)$ (1)
$\Delta W_{n,n}=W_{n,n}bone_{n,n}+bone_{n,n}$ (2)
Block Affine not only efficiently promotes information exchange between weights but also enhances the weight utilization of low-rank matrices. We conducted ablation experiments to verify the effectiveness of using a single learnable matrix structure. The comparison formulas are as follows:
$\Delta W_{n,n}=W_{n,n}bone\_A_{n,n}+bone\_B_{n,n}$ (3)
We know that in LLMs, the matrix $W$ is typically very large, and in most cases, the number of rows and columns are not equal. Setting $bone$ to the same dimensions as W would result in excessive computational overhead, which goes against the principles of fine-tuning. Therefore, we adjust $bone$ to be a low-rank matrix. When the weight matrix $W$ has dimensions $(n,m)$, it is divided according to a block_size $b$ into $W(n/b,b,m/b,b)$, yielding $(n/b)\times(m/b)$ blocks. Inspired by GQA(Grouped Query Attention)Ainslie et al. (2023), we group these blocks together to significantly reduce the number of trainable parameters without affecting performance. Different groups share a single Bone matrix for computation. Thus, the Bone matrix is set to $(g,b,b)$, where $g$ is the group size. The $W$ matrix is reshaped to $(gn,g,b,b)$, where $gn$ represents the number of groups. The Bone structure is illustrated in Figure 2(b). The final formula can be expressed as:
$W_{gn,g,b,b}=Reshape(W_{n,m})$ (4)
$\Delta W_{n,m}=Reshape(W_{gn,g,b,b}bone_{g,b,b}+bone_{g,b,b})$ (5)
During our comparative experiments between RWKV6 and LLaMA2, we found that the matrix $W$ has several different dimensions. For example, in RWKV6-7B, the dimension of ffn’s $k$ is $(4096*3.5,4096)$, while in LLaMA2-7B, the dimension of gate_proj in the mlp part is $(4096,11008)$. These differences require us to continuously adjust g to ensure the structure functions correctly. To better compare Bone with various LoRA variants and to address the compatibility issues arising from differences in FFN layers across different models and parameter settings, we modified the dimensions of Bone to $(n,b)$, where $b$ remains the size of the smallest block, and $n$ corresponds to the $n$ dimension of $W$. Thus, $g=n/b$. During computation, Bone only requires a dimension transformation, The Bone structure is illustrated in Figure 2(c), and the modified formula is as follows:
$W_{m/b,n/b,b,b}=Reshape(W_{n,m})$ (6)
$bone_{n/b,b,b}=Reshape(bone_{n,b})$ (7)
$\Delta W_{n,m}=Reshape(W_{m/b,n/b,b,b}bone_{n/b,b,b}+bone_{n/b,b,b})$ (8)
In this way, we can easily adapt to any dimension of W and models with different architectures. The compatibility method used above groups by columns, and naturally, we can also group by rows. The visualization is shown in Figure  2(d), and the formula is as follows:
$W_{n/b,m/b,b,b}=Reshape(W_{n,m})$ (9)
$bone_{m/b,b,b}=Reshape(bone_{m,b})$ (10)
$\Delta W_{n,m}=Reshape(W_{n/b,m/b,b,b}bone_{m/b,b,b}+bone_{m/b,b,b})$ (11)
(a) LoRA
(b) Bone-both
(c) Bone-col
(d) Bone-row
[…]
4 Experiments
The experiments were conducted on 4×NVIDIA 4090 24G GPUs.In our experiments, we fixed lora_r=32 to ensure consistent trainable parameters across different fine-tuning methods, using the AdamW optimizer with a batch size of 64, a learning rate of 2e-5 and a warmup ratio of 0.0, without any weight decay.
4.1 Evaluating the Performance of Bone under Different Parameter Settings in RWKV6
Table  1presents a comparative evaluation of the fine-tuned RWKV6 models with different parameter settings on the NLG task. I fine-tuned RWKV6 models with varying parameters on the MetaMathQA dataset Yu et al. (2023) and evaluated them on GSM8k Cobbe et al. (2021) and MATH datasets Yu et al. (2023). To ensure the accuracy of the experiment, the comparison between LoRA, PiSSA, and Bone was conducted with the same number of trainable parameters. The fine-tuning experiments for the 1.6B, 3B, and 7B RWKV6 models were each conducted for a complete epoch with a total batch size of 64 and a fixed learning rate of 2e-5.
Table 1: The performance of different PEFT techniques under various parameter settings on RWKV6.
Model Strategy Trainable GSM8K Math
RWKV 6-1.6B LoRA 33.0M 10.84 1.14
PiSSA 33.0M 12.89 1.28
Bone 33.0M 14.71 1.94
RWKV 6-3B LoRA 55.1M 21.22 2.34
PiSSA 55.1M 24.64 2.54
Bone 55.1M 25.25 2.82
RWKV 6-7B LoRA 88.1M 38.13 6.06
PiSSA 88.1M 40.48 6.12
Bone 88.1M 42.76 6.34
As shown in Table 1, Bone consistently outperforms PiSSA and LoRA across all parameter settings and tasks during fine-tuning. For example, on the GSM8k test, Bone achieved improvements of 1.82%, 0.61%, and 2.28% over PiSSA across different model sizes. This further demonstrates that the Bone architecture is adaptable to models with various parameter settings and consistently surpasses other fine-tuning structures.
4.2 Experiments on Llama2-7B
Models and Datasets
[…]
1. Math: We train our model on a 395k subset of MetaMathQA Yu et al. (2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K Cobbe et al. (2021) and MATH  Yu et al. (2023), with higher complexity and diversity.
2. Code: We train our model on a 100k subset of CodeFeedback Zheng et al. (2024b), a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval  Chen et al. (2021).
3. Chat: We train our model on a 70k subset of WizardLM-Evol-Instruct Xu et al. (2024). We test our model on the MT-Bench dataset  Zheng et al. (2024a), which consists of 80 multi-turn questions designed to assess LLMs on multiple aspects. The quality of the responses is judged by GPT-4o, and we report the first turn score.
Result
The results, as shown in Table  2 and Figure  1(a), demonstrate that Bone outperforms other structures in terms of convergence speed, data fitting, and generalization capabilities. Specifically, Bone outperforms PISSA by 5.84% on GSM8K and by 4.89% on HumanEval, further demonstrating the generalizability of the Bone approach.
Table 2: We fine-tuned LLaMA2-7B using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
Model Strategy Trainable GSM8K Math HumanEval MT-Bench
Llama2-7B LoRA 89.9M 38.97 0.44 16.46 3.93
OLoRA 89.9M 41.93 6.48 20.12 4.09
PiSSA 89.9M 43.52 6.92 20.73 4.25
Bone 87.0M 49.36 8.88 25.62 4.81
4.3 Experiments were conducted with different values of block_size $(b)$.
This subsection explores the upper limits of the Bone structure by varying the block_size in the Bone matrix. Notably, while we change the size of $b$, the value of $n$ remains constant. Therefore, as $b$ increases, the dimensions of Bone transform into the bone_g form, leading to a corresponding reduction in the number of blocks within each group. Comparative experiments were conducted by fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8K and Math benchmarks. The test results, as shown in Table  3, demonstrate that the fine-tuning performance improves as the value of b increases. Notably, when b reaches 16, the Bone structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark. However, its performance on the Math benchmark is weaker, indicating that the value of b plays a crucial role in determining the model’s generalization ability. As b increases, the model’s generalization capacity strengthens.
Table 3: Comparing different values of block_size $(b)$
Model Strategy Trainable GSM8K Math
Llama2-7B LoRA 89.9M 38.97 0.44
PiSSA 89.9M 43.52 6.92
Bone b=16 21.7M 45.79 3.54
Bone b=32 43.5M 46.70 7.67
Bone b=64 87.0M 49.36 8.88
Bone b=128 174.0M 53.15 10.56
Figure 3: Training loss curves of Bone with different block_size on the MetaMathQA dataset.
4.4 Ablation experiments on different grouping methods.
In this subsection, we explore the impact of different grouping methods in the Bone structure on model fine-tuning performance. Due to structural differences in the weight matrix W, the Bone-free grouping requires manual configuration, which is inconvenient. Therefore, this subsection only compares row-wise and column-wise grouping, both of which can be easily extended to any structure. We fine-tuned LLaMA2-7B on the MetaMathQA dataset and validated the results on GSM8k and Math. The results are shown in Table  4. Since in LLaMA2-7B, the dimension of gate_proj in the mlp part is $(4096,11008)$, this leads to an asymmetry between row-wise and column-wise grouping in the Bone structure, making it difficult to align parameter counts. Even though Bone-row performs better than Pissa with only around 20M parameters, the gap between Bone-row and Bone-col is too large. Initially, we suspected this gap was due to Bone-row lacking about 14M parameters compared to Bone-col. However, this assumption was disproved when we observed that Bone-row with 32 blocks and 43.5M parameters performed similarly to Bone-row with 72.8M parameters. This made it difficult to directly evaluate which grouping method performs better.
To explore the differences between the two grouping methods and the effect of parameter count, we added a comparative experiment with RWKV6-3B, as RWKV6’s symmetrical structure ensures that the trainable parameter count is the same whether using row-wise or column-wise grouping during Bone fine-tuning. This allows for a fairer comparison between Bone-row and Bone-col. The experimental results, shown in Table  4, indicate that the difference between the two is minimal, with both performing well. Therefore, we believe that Bone can effectively fit data regardless of the grouping method used. The key factor influencing Bone’s performance remains the block size. However, this doesn’t imply that different grouping methods are meaningless. As more LLMs begin to use techniques like GQA and MLA Liu et al. (2024a) to reduce KV cache overhead Dai et al. (2024); Lee et al. (2024); Shazeer (2019), the main weight matrices become smaller, and Bone will need to adjust its grouping or employ other techniques to adapt to these new technologies.
Table 4: Comparing Bone-row, Bone-col on math tasks
Model Strategy Trainable GSM8K Math
Llama2-7B Bone-row 72.8M 45.76 7.82
Bone-col 87.0M 49.36 8.88
RWKV6-3B Bone-row 44.1M 25.93 3.32
Bone-col 44.1M 25.25 2.82
4.5 Ablation experiments on Block Affine.
In this subsection, we aim to verify the effectiveness of a single learnable matrix, and we name the comparison structure Bone-AB. It consists of two low-rank matrices and is grouped by rows. To ensure the trainable parameter count is consistent, we set $b=32$ in Bone-AB and $b=64$ in Bone-row. For this comparison experiment, we continue to evaluate the Math task, with the results shown in Table  5. Bone-AB performed the worst, further demonstrating Block Affine not only efficiently promotes information exchange between weights but also enhances the weight utilization of low-rank matrices.
Table 5: Block Affine on math tasks
Model Strategy Trainable GSM8K Math
Llama2-7B Bone-col b=32 43.5M 46.70 7.67
Bone-row b=64 72.8M 45.76 7.82
Bone-AB b=32 72.8M 44.35 6.16
5 Resource and efficiency
Table  6 compares the training resources and token throughput required for fine-tuning RWKV6-3B using Bone and LoRA on a single 4090 GPU. The specific fine-tuning settings are as follows: batchsize = 1, context length (ctx_len) = 512. Due to the structure of Bone, an additional intermediate variable for the W weight is created during training, which increases memory usage by 5GB compared to LoRA. Additionally, Bone’s token throughput is only two-thirds of LoRA’s. To address this, we applied checkpointing specifically to the Bone structure, eliminating the extra intermediate variable, which successfully reduced memory usage but also further decreased token throughput.
Despite the fact that the Bone structure significantly outperforms LoRA variants in fine-tuning, we cannot ignore the efficiency issues. Therefore, a key focus of our future work will be improving the Bone operator to enhance token throughput and reduce memory usage.
Table 6: Resource and efficiency
Model Strategy Trainable GPU Memory Token throughput
RWKV-1.6B LoRA 33.0M 7314 MB 5.86 kt/s
Bone 33.0M 12276 MB 4.11 kt/s
Bone(checkpoint) 33.0M 7398 MB 3.46 kt/s
RWKV-3B LoRA 55.1M 12074 MB 3.62 kt/s
Bone 55.1M 22978 MB 2.16 kt/s
Bone(checkpoint) 55.1M 12168 MB 1.72 kt/s
6 Conclusion
This paper introduces a new PEFT technique called Bone(Block Affine), which employs block affine transformations on the weight matrix to capture the internal correlations within $W$, while also avoiding the complexity of initialization. Through extensive experiments, we found that Bone can achieve a rapid drop in loss during the early stages of training, similar to or even better than PISSA. Additionally, Bone demonstrates superior data fitting in the later stages of training and shows exceptional performance across different LLM architectures and parameter scales, proving its versatility. Bone brings new possibilities to existing LLM PEFT techniques. Instead of focusing solely on optimizing LoRA, we should shift our attention to innovative PEFT methods that are better suited to the architecture of LLMs.
7 Limitations and future work
Due to limited training resources, this paper lacks comparative experiments on full fine-tuning. However, the extensive experiments presented demonstrate that Bone possesses strong data fitting and generalization capabilities, paving the way for new directions in PEFT development. However, there are still some issues that require further consideration and resolution:
1. While the Bone architecture shows better training performance, its peak memory usage is higher compared to LoRA and its variants.
2. It remains to be seen whether Bone is suitable for large models in other modalities.
We welcome the community to provide additional suggestions and conduct further tests.