### v2

# Bone: Block Affine Transformation as Parameter Efficient Fine-tuning Methods for Large Language Models

Abstract
Low-Rank Adaptation (LoRA) has achieved remarkable training results by freezing the original weights and training only low-rank matrices, establishing itself as the predominant fine-tuning method for LLMs. In pursuit of performance closer to full-parameter training, a series of LoRA variants have emerged, such as LoRA+, PISSA, Olora, and LoRA-GA. However, these improvements complicate the initial setup of model training and increase initialization time. More importantly, they overlook the internal interactions of the original weight information. To address these issues, we introduce a novel theory, “Weight Guide” aimed at continuously guiding trainable matrices through the original weights during training to enhance the utilization of weight information. Based on this theory, we designed a new PEFT technique called Bone (Block Affine), which not only enhances the utilization of original weight information but also emphasizes the internal connections between weights, leading to faster convergence and better data fitting. Experimental comparisons across two different LLM architectures (LLaMA2, RWKV6) and various parameter scales demonstrate that the Bone structure can achieve rapid convergence and superior data fitting without the need for complex initialization. For example, when fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8k and math benchmarks, Bone achieved fine-tuning scores of 49.36 and 8.8, respectively, outperforming PISSA by 5.84% and 1.96%.
1 Introduction
Large models have been integrated into various industries, revolutionizing many traditional technologies Radford et al. (2019); Raffel et al. (2020). However, general-purpose large models often struggle to meet the needs of all downstream tasks, making it necessary to fine-tune base models for specific scenarios. Full-scale fine-tuning of large models is computationally costly; for example, finetuning the LLaMA2-7B Touvron et al. (2023) model with bfloat16 Wang & Kanwar (2019) precision requires around 60GB of VRAM. In contrast, PEFT (Parameter-Efficient Fine-Tuning)Xu et al. (2023) techniques could reduce the VRAM requirement to fit into a 24GB VRAM GPU. As a result, numerous PEFT techniques and quantization methods have emerged to reduce the training costs of large models. LoRA (Low-Rank Adaptation) Hu et al. (2021) has become one of the most popular PEFT methods due to its small tunable parameter size, its effectiveness, and the possibility of zero inference overhead after finetuning.
LoRA significantly reduces memory usage by freezing the original weights $W$ and updating two low-rank matrices $A$ and $B$. Typically, either $A$ or $B$ is initialized to zero, ensuring that the initial state of LoRA is consistent with the pre-trained model, The Figure 2(a) illustrates the structure visualization. However, extensive experiments Ding et al. (2023); Liu et al. (2024b); Biderman et al. (2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. This slow convergence is likely due to the small gradients caused by the zero initialization of either $A$ or $B$. To address this issue, researchers have proposed several LoRA variants, such as LoRA+ Hayou et al. (2024), PISSA Meng et al. (2024), and LoRA-GAWang et al. (2024). Despite their excellent performance, these LoRA variants inevitably introduce complexity into the fine-tuning process and reduce structural flexibility. For example, LoRA+ requires manual adjustment of different learning rates for $A$ and $B$; PISSA necessitates performing SVD decomposition on $W$ at the beginning of training, which can be time-consuming when the model parameters are large.
In this paper, we propose a novel PEFT method called Bone (Block Affine) to overcome the limitations of LoRA variants. It is well-known that LLMs have extremely large parameters, with complex interactions among their internal weights, but most LoRA variants utilize original weight information minimally, with only a few leveraging it during initialization to enhance performance. To increase the utilization of this information and promote internal interactions among weights, we propose a novel concept called “Weight Guide” This approach ensures that the trainable matrices are consistently guided and constrained by the original weights throughout every step of the training process. As illustrated in Figure  1(b), LoRA variants interact with the original weights only once during initialization to extract essential components. In contrast, the Bone structure is guided (constrained) by the original weights after each update, significantly increasing the new weights’ utilization of the original weights. As the training steps increase, the influence of the original weights becomes more pronounced. This is evident from the loss curves, which show that the differences between Bone and other methods grow larger in the later stages of training.
To promote the internal interactions of information within the original weight matrix $W$, Bone employs an trainable zero-initiated matrix, named bone, to perform Block Affine operations on $W$. The specific calculations of Block Affine are detailed in Section 3, and the structural visualization is shown in Figure  2. To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 Xu et al. (2023), RWKV6 Peng et al. (2024)). As shown in Figure  3, Bone achieves the fastest convergence, and the results on the test set demonstrate its superior data fitting and generalization capabilities (Table  2, Table  3).
Our contributions can be summarized as follows
1. We propose a new concept, “Weight Guide” which allows the continuous utilization of original weight information throughout the training process.
2. We propose a novel PEFT method called Bone (Block Affine) to overcome the limitations of LoRA variants. Block Affine not only efficiently promotes internal interactions of information within the weights but also enhances the weight utilization of low-rank matrices.
3. The effectiveness of the Bone structure and “Weight Guide” has been validated through extensive experiments.
(a) llama2-7b
(b) Train Steps
Figure 1: The left image shows the loss curve for LLaMA2-7B fine-tuned on the MetaMathQA dataset, with the first 100 steps highlighted for closer observation. Comparing the loss curves reveals that Bone demonstrates superior fitting ability across various architectures and parameter settings. Additionally, Bone exhibits a rapid decrease in loss within the first 100 steps, highlighting its effectiveness. The left image is a demonstration of the training state.
2 Related Works
The PEFT (Parameter-Efficient Fine-Tuning) techniques are diverse and include approaches like adapter tuningHoulsby et al. (2019); He et al. (2022); Wang et al. (2022); Pfeiffer et al. (2020), prefix tuningLiu et al. (2023); Li & Liang (2021), prompt tuningBrown (2020); Liu et al. (2023); Lester et al. (2021); Razdaibiedina et al. (2023); Li & Liang (2021), LoRAHu et al. (2021); Meng et al. (2024); Wang et al. (2024); Si et al. (2024), and layer-freezing methods such as LISA.
The adapter method does not require fine-tuning all the parameters of the pre-trained model. Instead, it introduces a small number of task-specific parameters to store knowledge related to that task, thereby reducing the computational demands of model fine-tuning. Prefix tuning is a lightweight fine-tuning method for generative tasks. It adds a continuous, task-specific vector sequence, called a prefix, to the input. Unlike prompts, prefixes are entirely composed of free parameters and do not correspond to actual tokens. Compared to traditional fine-tuning, prefix tuning only optimizes a set of prefixes related to specific tasks, without any change to the original model.
Prompt tuning defines a unique prompt for each task, prepending it to the input data while freezing the pre-trained model during training. LoRA allows us to indirectly train certain dense layers in a neural network by optimizing low-rank decomposition matrices that adapt these layers while keeping the pre-trained weights unchanged. This approach significantly addresses the inference latency introduced by adapters and the fitting challenges of prefix tuning.
In LoRA, the adapter matrices $A$ and $B$ are updated with the same learning rate, but using the same rate for both may not effectively learn the features. LoRA+ extended this method by introducing independent learning rates for matrices $A$ and $B$ with a fixed ratio, improving the method’s efficiency. The DoRA Liu et al. (2024b) method combines weight decomposition to achieve learning capabilities similar to full fine-tuning without sacrificing LoRA’s inference efficiency. PiSSA optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), PiSSA initializes the dominant singular values and vectors to train these matrices, while keeping the residual matrix static during fine-tuning. OLoRA Büyükakyüz (2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $A$ and $B$ by computing the initial gradient, thereby closely approximating full fine-tuning.
3 Bone: Block Affine
In this section, we will introduce the Bone (Block Affine) method. Our design is based on multiple considerations. First, to achieve the aforementioned “Weight Guide” We modify the formula $Y=X(W+\Delta W)$ by replacing $\Delta W$ with $W\Delta W+\Delta W$, allowing the original weights to continuously guide (constrain) the direction of $\Delta W$ throughout the training process. To verify whether the guidance (constraint) from the original weights has a positive effect on the trainable matrix, we conducted an ablation experiment in Section  4.5, we named the comparison structure “Bone-Unconstrained”, with the comparison formula: $\Delta W=WA+B$. The experimental results indicate that “Weight Guide” not only provides positive guidance but also enhances the utilization of matrix parameters.
Second, to enhance the fusion of internal weight features and extract more effective information, we opted for matrix multiplication over the Hadamard product in the $W\Delta W$ operation. The Hadamard product only captures features at corresponding positions, whereas matrix multiplication allows the accumulation of features from an entire row into a single feature. To validate this approach, we conducted experiments in Section  4.5 to compare different computational methods in terms of feature fusion and extraction. Here, we named the comparison structure “Bone-Hadamard” to facilitate readers’ understanding and reference in the following sections. Ultimately, we chose matrix multiplication as the core operation.
Finally, we hypothesize that the matrix $W$ has dimensions $(n,n)$, ideally, we train with a full-rank matrix $bone$ $(n,n)$. The specific calculation for Block Affine is as follows:
$Y=X(W+\Delta W)$ (1)
$\Delta W_{n,n}=W_{n,n}bone_{n,n}+bone_{n,n}$ (2)
We know that in LLMs, the matrix $W$ is typically very large, and in most cases, the number of rows and columns is not equal. Setting $bone$ to the same dimensions as W would result in excessive computational overhead, which goes against the principle of parameter-efficiency. Therefore, we adjust $bone$ to be a low-rank matrix. When the weight matrix $W$ has dimensions $(n,m)$, it is divided according to a block_size $b$ into $W(n/b,b,m/b,b)$, yielding $(n/b)\times(m/b)$ blocks. Inspired by GQA(Grouped Query Attention)Ainslie et al. (2023), we group these blocks together to significantly reduce the number of trainable parameters without affecting performance. Different groups share a single Bone matrix for computation. Thus, the Bone matrix is set to $(g,b,b)$, where $g$ is the group size. The $W$ matrix is reshaped to $(gn,g,b,b)$, where $gn$ represents the number of groups. The Bone structure is illustrated in Figure 2(b). The final formula can be expressed as:
$W_{gn,g,b,b}=Reshape(W_{gn*b,g*b})=Reshape(W_{n,m})$ (3)
$\Delta W_{n,m}=Reshape(W_{gn,g,b,b}bone_{g,b,b}+bone_{g,b,b})$ (4)
During our comparative experiments between RWKV6 and LLaMA2, we found that the matrix $W$ has several different dimensions. For example, in RWKV6-7B, the dimension of ffn’s $k$ is $(4096*3.5,4096)$, while in LLaMA2-7B, the dimension of gate_proj in the mlp layer is $(4096,11008)$. These differences require us to continuously adjust g to ensure the structure functions correctly. To better compare Bone with various LoRA variants and to address the compatibility issues arising from differences in FFN layers across different models and parameter settings, We set $g=n/b$ in $bone_{g,b,b}$ allowing the parameters of Bone to be equivalent to those of $LoraA_{n,r}$. During computation, Bone only requires a dimension transformation, The Bone structure is illustrated in Figure 2(c), and the modified formula is as follows:
$W_{m/b,n/b,b,b}=Reshape(W_{n,m})$ (5)
$bone_{n/b,b,b}=Reshape(bone_{n,b})$ (6)
$\Delta W_{n,m}=Reshape(W_{m/b,n/b,b,b}bone_{n/b,b,b}+bone_{n/b,b,b})$ (7)
In this way, we can easily adapt to any dimension of W and models with different architectures. The compatibility method used above groups by columns, and naturally, we can also group by rows. The visualization is shown in Figure  2(d), and the formula is as follows:
$W_{n/b,m/b,b,b}=Reshape(W_{n,m})$ (8)
$bone_{m/b,b,b}=Reshape(bone_{m,b})$ (9)
$\Delta W_{n,m}=Reshape(W_{n/b,m/b,b,b}bone_{m/b,b,b}+bone_{m/b,b,b})$ (10)
(a) LoRA
(b) Bone-both
(c) Bone-col
(d) Bone-row
[…]
4 Experiments
The experiments were conducted on 4×NVIDIA 4090 24G GPUs. Our experimental training parameter settings are shown in the table  1.
Table 1: Training parameter settings
Deivce Model low rank gobal batchsize learning rate precision
4×NVIDIA 4090 RWKV6 r=32/b=64 64 2e-5 bf16
Llama r=36/b=64 64 2e-5 bf16
4.1 Evaluating the Performance of Bone under Different Parameter Settings in RWKV6
Models and Datasets
To validate the stability of Bone, we trained RWKV6 models with different parameter sizes on math tasks.
Math: We fine-tuned RWKV6 models with varying parameters on the MetaMathQA dataset Yu et al. (2023) and evaluated them on GSM8k Cobbe et al. (2021) and MATH datasets Yu et al. (2023).
Implementation Details
For fine-tuning RWKV6 models with different parameter sizes, we kept the learning rate, batch size, and training steps the same across all experiments. We set lora_r=32 and bone_b=64 to ensure that the trainable parameters are equivalent between the two methods.
Result
As shown in Table 2, Bone consistently outperforms PiSSA and LoRA across all parameter settings and tasks during fine-tuning. For example, on the GSM8k test, Bone achieved improvements of 1.82%, 0.61%, and 2.28% over PiSSA across different model sizes. This further demonstrates that the Bone architecture is adaptable to models with various parameter settings and consistently surpasses other fine-tuning structures.
(a) RWKV6 loss
Figure 3: The image displays the loss curves for RWKV6 with different parameter settings, fine-tuned on MetaMathQA.
Table 2: The performance of different PEFT techniques under various parameter settings on RWKV6.
Model Strategy Trainable GSM8K Math
RWKV 6-1.6B LoRA 33.0M 10.84 1.14
PiSSA 33.0M 12.89 1.28
Bone 33.0M 14.71 1.94
RWKV 6-3B LoRA 55.1M 21.22 2.34
PiSSA 55.1M 24.64 2.54
Bone 55.1M 25.25 2.82
RWKV 6-7B LoRA 88.1M 38.13 6.06
PiSSA 88.1M 40.48 6.12
Bone 88.1M 42.76 6.34
4.2 Experiments on Llama2-7B
Models and Datasets
To verify the generalizability of Bone, we conducted more comprehensive experiments on LLaMA2-7B. we conducted 3 more task finetuning experiments on LLaMa2-7B: math, code, and chat.
1. Math: We trained our model on a 395k subset of MetaMathQA Yu et al. (2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K Cobbe et al. (2021) and MATH  Yu et al. (2023), with higher complexity and diversity.
2. Code: We train our model on a 100k subset of CodeFeedback Zheng et al. (2024b), a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval  Chen et al. (2021).
3. Chat: We train our model on a 70k subset of WizardLM-Evol-Instruct Xu et al. (2024). We test our model on the MT-Bench dataset  Zheng et al. (2024a), which consists of 80 multi-turn questions designed to assess LLMs on multiple aspects. We used GPT-4o to judge the quality of responses, as shown in lm-sys/FastChat.
Implementation Details
[…]
Result
The results, as shown in Table  3 and Figure  1(a), demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities. Specifically, Bone outperforms PISSA by 5.84% on GSM8K and by 4.89% on HumanEval, further Bone outperformed methods compared, on all 3 different tasks focusing on various disciplines, demonstrating the generalizability.
Table 3: We fine-tuned LLaMA2-7B using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
Model Strategy Trainable GSM8K Math HumanEval MT-Bench
Llama2-7B LoRA 89.9M 38.97 0.44 16.46 3.93
OLoRA 89.9M 41.93 6.48 20.12 4.09
PiSSA 89.9M 43.52 6.92 20.73 4.25
Bone 87.0M 49.36 8.88 25.62 4.81
4.3 Experiments on different values of block_size $(b)$.
This subsection explores the upper limits of the Bone structure by varying the block_size in the Bone matrix. Notably, while we change the size of $b$, the value of $n$ remains constant. Therefore, as $b$ increases, the dimensions of Bone transform into the bone_g form, leading to a corresponding reduction in the number of blocks within each group. Comparative experiments were conducted by fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8K and Math benchmarks. The test results, as shown in Table  4, demonstrate that the fine-tuning performance improves as the value of b increases. Notably, when b reaches 16, the Bone structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark. However, its performance on the Math benchmark is only 3.54. The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of b impacts the model’s ability to understand unseen data. Based on this observation, we conclude that increasing b not only raises the number of parameters but also enlarges the window for information fusion, allowing the model to perform better on unseen data.
Table 4: Comparing different values of block_size $(b)$
Model Strategy Trainable GSM8K Math
Llama2-7B LoRA 89.9M 38.97 0.44
PiSSA 89.9M 43.52 6.92
Bone b=16 21.7M 45.79 3.54
Bone b=32 43.5M 46.70 7.67
Bone b=64 87.0M 49.36 8.88
Bone b=128 174.0M 53.15 10.56
Figure 4: Training loss curves of Bone with different block_size on the MetaMathQA dataset.
4.4 Ablation experiments on different grouping methods.
In this subsection, we explore the impact of different grouping methods in the Bone structure on model fine-tuning performance. Due to structural differences in the weight matrix W, the Bone-free grouping requires manual configuration, which is inconvenient. Therefore, this subsection only compares row-wise and column-wise grouping, both of which can be easily extended to any structure. We fine-tuned LLaMA2-7B on the MetaMathQA dataset and validated the results on GSM8k and Math. The results are shown in Table  5. Since in LLaMA2-7B, the dimension of gate_proj in the mlp part is $(4096,11008)$, this leads to an asymmetry between row-wise and column-wise grouping in the Bone structure, making it difficult to align parameter counts. Even though Bone-row performs better than Pissa with only around 20M parameters, the gap between Bone-row and Bone-col is too large. Initially, we suspected this gap was due to Bone-row lacking about 14M parameters compared to Bone-col. However, this assumption was disproved when we observed that Bone-row with 32 blocks and 43.5M parameters performed similarly to Bone-row with 72.8M parameters. This made it difficult to directly evaluate which grouping method performs better.
To explore the differences between the two grouping methods and the effect of parameter count, we added a comparative experiment with RWKV6-3B, as RWKV6’s symmetrical structure ensures that the trainable parameter count is the same whether using row-wise or column-wise grouping during Bone fine-tuning. This allows for a fairer comparison between Bone-row and Bone-col. The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well. Therefore, we believe that Bone can effectively fit data regardless of the grouping method used. The key factor influencing Bone’s performance remains the block size. However, this doesn’t imply that different grouping methods are meaningless. As more LLMs begin to use techniques like GQA and MLA Liu et al. (2024a) to reduce KV cache overhead Dai et al. (2024); Lee et al. (2024); Shazeer (2019), the main weight matrices become smaller, and Bone will need to adjust its grouping or employ other techniques to adapt to these new technologies.
Table 5: Comparing Bone-row, Bone-col on math tasks
Model Strategy Trainable GSM8K Math
Llama2-7B Bone-row 72.8M 45.76 7.82
Bone-col 87.0M 49.36 8.88
RWKV6-3B Bone-row 55.1M 25.93 3.32
Bone-col 55.1M 25.25 2.82
4.5 Ablation experiments on Block Affine.
We will conduct ablation experiments on the Block Affine structure, which consist of two parts:
1. Verifying the effectiveness of “Weight Guide” on LLaMA2-7B: The results are shown in the table  6. It is evident that the performance of Bone-Unconstrained is even worse than that of Bone-Col $b=32$, which has half the number of parameters but incorporates the “Weight Guide” characteristic.
2. Evaluating the best computation method for promoting internal weight feature fusion on RWKV6-3B: As seen in the table  6, Bone-Hadamard is significantly weaker compared to matrix multiplication.
The comparison formulas are as follows:
$Bone-Unconstrained\Delta W_{n,n}=W_{n,n}A_{n,n}+B_{n,n}$ (11)
$Bone-Hadamard\Delta W_{n,n}=W_{n,n}*bone_{n,n}+bone_{n,n}$ (12)
Table 6: Block Affine on math tasks
Model Strategy Trainable GSM8K Math
Llama2-7B Bone-col b=32 43.5M 46.70 7.67
Bone-row b=64 72.8M 45.76 7.82
Bone-Unconstrained b=32 72.8M 44.35 6.16
RWKV6-3B Bone 55.1M 25.93 3.32
Bone-Hadamard 55.1M 22.44 2.67
5 Resource and efficiency
Table  7 compares the training resources and token throughput required for fine-tuning RWKV6-3B using Bone and LoRA on a single 4090 GPU. The specific fine-tuning settings are as follows: batchsize = 1, context length (ctx_len) = 512. The structure of Bone, an additional intermediate variable for the W weight is created during training, which increases memory usage. To address this, we applied checkpointing specifically to the Bone structure, eliminating the extra intermediate variable, which successfully reduced memory usage but also further decreased token throughput. At the end of the table, we provide the real-world training scenarios comparing the memory usage and token throughput of Bone and LoRA.
Therefore, a key focus of our future work will be improving the Bone operator to enhance token throughput and reduce memory usage.
Table 7: Resource and efficiency
Model Strategy Trainable GPU Memory Token throughput
RWKV-1.6B LoRA 33.0M 7314 MB 5.86 kt/s
Bone 33.0M 12276 MB 4.11 kt/s
Bone(checkpoint) 33.0M 7398 MB 3.46 kt/s
Bone(model checkpoint) 33.0M 7398 MB 1.59 kt/s
RWKV-3B LoRA 55.1M 12074 MB 3.62 kt/s
Bone 55.1M 22978 MB 2.16 kt/s
Bone(checkpoint) 55.1M 12168 MB 1.72 kt/s
RWKV-3B LoRA(model checkpoint) 55.1M 4*15328 MB 15.6 kt/s
Bone(model checkpoint) 55.1M 4*15305 MB 14.2 kt/s
6 Conclusion
This paper introduces a novel theory, “Weight Guide” aimed at continuously guiding trainable matrices through the original weights during training to enhance the utilization of weight information. Based on this theory, we designed a new PEFT technique called Bone (Block Affine). Additionally, we employed block matrix multiplication to efficiently promote information fusion within the weights in the Bone structure. Through extensive experiments, we found that Bone can achieve a rapid drop in loss during the early stages of training, similar to or even better than PISSA. Additionally, Bone demonstrates superior data fitting in the later stages of training and shows exceptional performance across different LLM architectures and parameter scales, proving its versatility. Bone brings new possibilities to existing LLM PEFT techniques. Instead of focusing solely on optimizing LoRA, we should shift our attention to innovative PEFT methods that are better suited to the architecture of LLMs.
7 Limitations and future work
Due to limited training resources, this paper lacks comparative experiments on full fine-tuning. However, the extensive experiments presented demonstrate that Bone possesses strong data fitting and generalization capabilities, paving the way for new directions in PEFT development. However, there are still some issues that require further consideration and resolution:
1. While the Bone architecture shows better training performance, its peak memory usage is higher compared to LoRA and its variants.
2. It remains to be seen whether Bone is suitable for large models in other modalities.
We welcome the community to provide additional suggestions and conduct further tests.