### v6

# DiSHA: Dimension-Sharding Adaptation with Fast Convergence and Fast Computation

Abstract
Low-Rank Adaptation (LoRA) leverages the low intrinsic rank of weight updates in Large Language Models (LLMs), establishing a Parameter-Efficient Fine-Tuning (PEFT) paradigm. However, LoRA suffers from slow convergence. We introduce Dimension-Sharding Adaptation (DiSHA), which expands the PEFT design space to unlock lower intrinsic ranks and faster convergence by default. Within DiSHA’s design space, we propose Block Affine Adaptation (Bone), a computationally efficient structure that delivers both high performance and efficiency. While certain DiSHA configurations may result in colinear updates to weight shards, we address this with Block Affine Transformation Adaptation (BAT), a nonlinear variant of DiSHA. BAT introduces nonlinearity by combining trainable matrices with original weight shards in a nonlinear manner, inducing nonlinearity in matrix updates without introducing additional parameters. Empirical results show that Bone, under the DiSHA framework, consistently outperforms LoRA variants in both NLG and NLU tasks, with significantly improved computational efficiency. Further analysis demonstrates that BAT enhances model capabilities by leveraging its nonlinear design.
1 Introduction
The emergence of Large Language Models (LLMs) has fundamentally transformed many traditional technologies  Radford et al. (2019); Raffel et al. (2020). However, general-purpose large models often struggle to meet the needs of all downstream tasks, making it necessary to fine-tune base models for specific scenarios. Full-scale fine-tuning of large models is computationally costly. As a result, numerous Parameter-Efficient Fine-Tuning (PEFT)Xu et al. (2023) techniques and quantization methods have emerged to reduce the training costs of large models. Low-Rank Adaptation(LoRA) Hu et al. (2021) has become one of the most popular PEFT methods due to its small tunable parameter size, its effectiveness, and the possibility of zero inference overhead after fine-tuning. The figure  3 illustrates the structure visualization.
32641282560612182430rankgradient normftfrozen ftLoRABigLoRAPiSSADiSHA
Figure 1: the initial gradient norm of model training.
However, extensive experiments Ding et al. (2023); Liu et al. (2024b); Biderman et al. (2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. To address this issue, researchers have proposed several LoRA variants, such as LoRA+ Hayou et al. (2024), PISSA Meng et al. (2024), and LoRA-GAWang et al. (2024). These methods have all brought unexpected improvements. By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed.
Different initializations of LoRA variants accelerate convergence essentially by increasing the initial gradients during training or aligning them with the full-scale training gradients. However, this also makes the initialization process significantly more complex, leading to increased initialization time overhead. To address the issue of complexity and explore the specific reasons for LoRA’s slow convergence, we conducted extensive experiments on initial gradient testing. Ultimately, we concluded that LoRA’s slow convergence is influenced not only by initialization but also by its inherent structural limitations. A simpler update structure may be more suitable for fine-tuning LLMs.
In this work, we first propose a novel framework called Dimension-Sharding Adaptation (DiSHA) to design PEFT with lower ranks and faster convergence. In DiSHA, the pre-trained weights $W_{0}$ are frozen and divided into multiple shards during training, with all shard weights sharing a trainable matrix of the same size as a shard to perform updates, significantly reducing the number of trainable parameters. The design space of DiSHA is vast, allowing pre-trained weights to be divided into dimensions of arbitrary sizes 2. Computational efficiencies vary among design choices. We proposed an efficient and user-friendly structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments. Of course, we also observed that certain subsets of the DiSHA design space exhibit limitations: all shard weight updates are colinear, which could potentially hinder model performance. To address this potential issue, we propose the Bat method. Without introducing additional trainable matrices, Bat uses the trainable matrix to interact with the original weight shards, transforming the linear updates between the shards into non-linear updates. This approach enhances the model’s capability while maintaining efficiency.
Figure 2: partition of weight matrix $W_{0}$.
Our extensive evaluation shows that Bone demonstrates excellent performance on both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks, significantly surpassing LoRA and its variants. Moreover, as shown in Figure  5, Bone also achieves a significantly faster convergence rate compared to PiSSA. This demonstrates the feasibility of the DiSHA framework. Furthermore, Bone’s unique design not only resolves the unification challenges of the DiSHA framework, allowing seamless adaptation across different LLM architectures such as LLaMA, RWKV, and Mistral but also enables it to significantly outperform LoRA in both computational efficiency and memory usage. In NLG tasks, Bone demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variant Pissa across the board. To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 Xu et al. (2023), RWKV6 Peng et al. (2024)). Based on the comparative experiments between Bone and Bat, it is evident that the non-linear updates of weight shards enhance the model’s performance ceiling.
Our contributions can be summarized as follows
1. We propose a novel PEFT framework called DiSHA which significantly reduces the number of trainable parameters.
2. We proposed an efficient and user-friendly structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments.
3. We identified a potential issue in the DiSHA framework: the updates of weight shards are colinear. To address this, we improved upon Bone and proposed Bat, which transforms weight shard updates into non-linear ones, thereby enhancing the model’s performance ceiling.
4. Through extensive experiments, we validated the effectiveness of DiSHA, Bone, and Bat, achieving state-of-the-art (SOTA) performance.
2 Related Works
The PEFT (Parameter-Efficient Fine-Tuning) techniques are diverse and include approaches like adapter tuningHoulsby et al. (2019); He et al. (2022); Wang et al. (2022); Pfeiffer et al. (2020), prefix tuningLiu et al. (2023); Li & Liang (2021), prompt tuningBrown (2020); Liu et al. (2023); Lester et al. (2021); Razdaibiedina et al. (2023); Li & Liang (2021), LoRAHu et al. (2021); Meng et al. (2024); Wang et al. (2024); Si et al. (2024), and layer-freezing methods such as LISA.
LoRA, as one of the most popular PEFT techniques, has demonstrated that the updates to the weights exhibit a low intrinsic rank during adaptation. In LoRA, for a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. During training, $W_{0}$ is frozen and does not receive gradient updates, while $A$ and $B$ contain trainable parameters. Note both $W_{0}$ and $\Delta W=BA$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise. For $h=W_{0}x$, our modified forward pass yields:
$h=W_{0}x+\Delta Wx=W_{0}x+BAx$ (1)
Since A and B are initialized with Gaussian noise and zeros in LoRA, the gradients can be very small, leading to slow convergence in the fine-tuning process. As a result, many LoRA variants have emerged, aiming to enhance LoRA’s capabilities by modifying the initialization or initial gradients, among other approaches.
LoRA+ extended this method by introducing independent learning rates for matrices $A$ and $B$ with a fixed ratio, improving the method’s efficiency. The DoRA Liu et al. (2024b) method combines weight decomposition to achieve learning capabilities similar to full fine-tuning without sacrificing LoRA’s inference efficiency. PiSSA optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), PiSSA initializes the dominant singular values and vectors to train these matrices, while keeping the residual matrix static during fine-tuning. OLoRA Büyükakyüz (2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $A$ and $B$ by computing the initial gradient, thereby closely approximating full fine-tuning.
The improvements brought by these LoRA variants validate that the updates to the weights exhibit a low intrinsic rank during adaptation and hold greater potential. However, they also introduce more complex initialization steps and increase preprocessing time. Therefore, we need to design novel structures to facilitate the exploration of a lower intrinsic rank.
3 Method
3.1 DiSHA: Dimension-Sharding Adaptation
Hypothesis Based on the conclusions of the experimental analysis of the gradient norm  9.1, we designed a novel framework called Dimension-Sharding Adaptation (DiSHA) to unlock lower intrinsic ranks and faster convergence by default. For a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $W=W_{0}+\Delta W=W_{0}+expand(D)$, where $D\in\mathbb{R}^{r_{1}\times r_{2}},expand(D)\in\mathbb{R}^{d\times k}$, and the rank $(r_{1},r_{2})\ll\min(d,k)$. During training, $W_{0}$ is frozen and does not receive gradient updates, while $D$ is a trainable parameter initialized to 0. Note both $W_{0}$ and $\Delta W=expand(D)$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise. For $y=W_{0}x$, where $x\in\mathbb{R}^{b\times l\times k}$, $y\in\mathbb{R}^{b\times l\times d}$, our modified forward pass yields:
$y=W_{0}x+\Delta Wx=W_{0}x+expand(D)x$ (2)
Here, we explain the concept of $expand(D)$. As shown in the DiSHA structure in Figure  3, we divide $W_{0}$ into multiple shards along the dimension of $D$. The $D$ matrix is then replicated according to the number of shards, and its dimensions are restored to match $W_{0}$. Finally, the updated shards are summed to update $W$.
3.2 DiSHA design space
In the previous section, we introduced the design principles of DiSHA and defined its trainable matrix $D\in\mathbb{R}^{r_{1}\times r_{2}}$. With DiSHA, you can customize the size of the trainable matrix $D$ based on the dimensions of the pre-trained weights. Since all shards of the pre-trained weights $W_{0}$ share a single $D$ for updates, as long as the expanded $expand(D)$ matches the dimensions of $W_{0}$, you can freely choose how to divide $W_{0}$ into shards. This flexibility makes the design space of DiSHA immense 2.
3.3 Bone: Block-Affine efficient update
In the previous section, we introduced the DiSHA framework, a method that significantly reduces the number of trainable parameters in the model. Due to the vast design space of the DiSHA framework, it is difficult for users to accurately determine suitable values for training. A good setting of $D\in\mathbb{R}^{r_{1}\times r_{2}}$ does not generalize well across all neural networks. For example, with the rise of technologies like GQA Ainslie et al. (2023) and MQA Shazeer (2019), the matrix dimensions of self-attention modules ($W_{k},W_{v}$) vary greatly among different LLM models. At the same time, computational efficiency and resource utilization also vary depending on the design space choices within the DiSHA framework.
To address these limitations, we proposed Block Affine Adaptation (Bone) based on the DiSHA framework. Inspired by LoRA, we modified $D\in\mathbb{R}^{r_{1}\times r_{2}}$ to $D\in\mathbb{R}^{r\times d}$ or $D\in\mathbb{R}^{r\times k}$. This allows the trainable parameter count to be controlled by a single parameter $r$, making it easy to adapt to any neural network. Additionally, we observed that $expand(D)x$ not only increases memory consumption but also reduces computational efficiency. From our analysis, when $D\in\mathbb{R}^{r\times d}$, $x\in\mathbb{R}^{b\times l\times k}$, $expand(D)\in\mathbb{R}^{d\times k}$, the operation $expand(D)x$ essentially performs a summation along the $k$-dimension, where $expand(D)$ involves replicating the necessary number of shards. Clearly, we can first $sum(x)$ along the $k$dimension and then compute it directly with $D$, where $sum(x)\in\mathbb{R}^{b\times l\times r}$. This approach not only reduces memory consumption but also improves computational efficiency. The update for $W$ is shown in the figure  3, and the Bone calculation formula is as follows:
$y=W_{0}x+\Delta Wx=W_{0}x+sum(x)D$ (3)
Figure 3: Diagram of LoRA, DiSHA, and Bone Architectures.
3.4 Bat: Block-Affine Transformation
We conducted extensive experiments on both NLU and NLG tasks to validate the effectiveness of Bone. It outperforms many LoRA variants and surpasses LoRA in terms of memory consumption and computational efficiency. However, Bone still has a potential issue, which stems from the shared update mechanism of the DiSHA framework. Specifically, all shards in the weights use the same trainable matrix for updates, causing the updates of all shards to be colinear. This restricts weight updates and, consequently, the model’s capabilities. To address the issue of linear correlation among shard updates, our initial idea was to use a trainable coefficient matrix to control the updates of different shards. However, this approach would increase additional parameters.
To avoid introducing additional parameters, we applied a linear transformation between the original weight $W_{0}$ shards and the trainable matrix, ensuring that the updates of different shards are nonlinearly related. Based on this, we proposed the Block Affine Transformation Adaptation (Bat) method. It extends Bone by further dividing both the trainable matrix and $W_{0}$ into smaller shards, where each shard interacts with the trainable matrix $D$ through a linear transformation, achieving nonlinear independence among shard updates. The specific formula is as follows:
$W_{k/r,d/r,r,r}$ $=Reshape(W_{d,k})$ (4)
$bone_{d/r,r,r}$ $=Reshape(bone_{r,d})$
$\Delta W_{d,k}$ $=Reshape(W_{k/r,d/r,r,r}bone_{d/r,r,r}+bone_{d/r,r,r})$
It is evident that Bat can be defined in many other ways, as illustrated in the figure 4. In Section 6, we compared the performance of different partitioning strategies for the structure. The formula for Bone-row is as follows:
$W_{d/r,k/r,r,r}$ $=Reshape(W_{d,k})$ (5)
$bone_{k/r,r,r}$ $=Reshape(bone_{r,k})$
$\Delta W_{d,k}$ $=Reshape(W_{d/r,k/r,r,r}bone_{k/r,r,r}+bone_{k/r,r,r})$
Figure 4: Diagram of Bat Structures in Different Design Spaces.
4 Experiments
In this section, we evaluate the performance of Bone on various benchmark datasets. Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset with the robert-base model. Subsequently, we evaluated the Natural Language Generation (NLG) capabilities by fine-tuning the LLM.
The experiments were conducted on 4×NVIDIA 4090 24G GPUs.
4.1 Experiments on Natural Language Understanding
Models and Datasets
We fine-tune the RoBERTa-base model on several datasets from the GLUE benchmark, including MNLI, SST-2, CoLA, QNLI, and MRPC. Performance is evaluated on the development set using accuracy as the primary metric.
Implementation Details
The experimental hyperparameter settings were aligned with those in the LoRA repository, but training was conducted using a single 4090 GPU. Each experiment is conducted with 3 different random seeds, and the average performance is reported.
Results
As shown in Table 1, Bone demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and Pissa.
Table 1: The results of fine-tuning RoBERTa-base using Bone and various LoRA variants were compared on a subset of the GLUE benchmark.
Method Trainable MNLI SST-2 CoLA QNLI MRPC
LoRA 0.236% $85.63_{\pm 0.01}$ $\mathbf{94.03_{\pm 0.02}}$ $62.40_{\pm 0.71}$ $91.37_{\pm 0.97}$ $87.98_{\pm 0.23}$
Pissa 0.236% $\mathbf{85.72_{\pm 0.40}}$ $93.64_{\pm 0.13}$ $67.28_{\pm 0.59}$ $91.40_{\pm 0.54}$ $88.11_{\pm 0.24}$
Bone 0.236% $85.71_{\pm 0.32}$ $93.60_{\pm 0.07}$ $\mathbf{72.86_{\pm 3.13}}$ $\mathbf{91.43_{\pm 0.76}}$ $\mathbf{88.14_{\pm 0.60}}$
4.2 Experiment on Natural Language Generation
Models and Datasets
To verify the generalizability of Bone, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math, code, and chat.
1. Math: We trained our model on a 395k subset of MetaMathQA Yu et al. (2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K Cobbe et al. (2021) and MATH  Yu et al. (2023), with higher complexity and diversity.
2. Code: We train our model on a 100k subset of CodeFeedback Zheng et al. (2024b), a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval  Chen et al. (2021).
3. Chat: We train our model on a 70k subset of WizardLM-Evol-Instruct Xu et al. (2024). We test our model on the MT-Bench dataset  Zheng et al. (2024a), which consists of 80 multi-turn questions designed to assess LLMs on multiple aspects. We used GPT-4o to judge the quality of responses, as shown in lm-sys/FastChat.
Figure 5: The image shows the loss curve for LLaMA2-7B fine-tuned on the MetaMathQA dataset, with the first 100 steps highlighted for closer observation. Comparing the loss curves reveals that Bone demonstrates superior fitting ability across various architectures and parameter settings. Additionally, Bone exhibits a rapid decrease in loss within the first 100 steps, highlighting its effectiveness.
Implementation Details
The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used. It is worth noting that the weights of LLaMA2-7B are not fully symmetric, making it impossible to perfectly align the trainable parameters when comparing Bone and LoRA. To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of Bone to 64, ensuring that Bone uses fewer parameters than LoRA to demonstrate its superiority. Each experiment is conducted with 2 different random seeds, and the average performance is reported.
Result
The results, as shown in Table  2 and Figure  5, demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities. Bone demonstrates outstanding performance across three different tasks. On LLaMA2-7B, Bone achieves results that surpass Pissa, despite using fewer parameters than LoRA and its variants. On RWKV6-7B, Bone and LoRA have the same number of trainable parameters, yet Bone consistently outperforms LoRA and its variants across all tasks.
Table 2: We fine-tuned LLMs using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
Model Strategy Trainable GSM8K Math HumanEval MT-Bench
Llama2-7B LoRA 89.9M 40.75 5.22 17.68 3.73
OLoRA 89.9M 42.93 6.51 21.12 4.03
PiSSA 89.9M 43.89 6.92 22.25 4.11
Bone 87.0M 48.16 8.58 24.08 4.31
RWKV 6-7B LoRA 88.1M 38.13 6.06 - -
PiSSA 88.1M 40.48 6.12 - -
Bone 88.1M 41.73 6.52 - -
Mistral-7B LoRA 89.1M 65.17 15.82 39.02 -
PiSSA 89.1M 67.01 18.13 40.85 -
Bone 88.1M 66.94 18.85 41.76 -
4.3 Effect of Rank $r$
This subsection explores the upper limits of the Bone structure by varying the rank $r$ in the Bone matrix. Comparative experiments were conducted by fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8K and Math benchmarks. The test results, as shown in Table  3, demonstrate that the fine-tuning performance improves as the value of b increases. Notably, when $r=16$, the Bone structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark. However, its performance on the Math benchmark is only 3.73. The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of $r$ impacts the model’s ability to understand unseen data. Based on this observation, we hypothesize that when the rank is too small, it significantly limits the model’s generalization ability.
Table 3: Comparing different values of rank $(r)$
Model rank Trainable GSM8K Math
Llama2-7B 16 21.7M 45.90 3.77
32 43.5M 46.18 7.43
64 87.0M 48.16 8.58
128 174.0M 53.49 10.08
Figure 6: Training loss curves of Bone with different rank $r$ on the MetaMathQA dataset.
4.4 Bone vs Bat
Table 4: Comparing Bone, Bat on math tasks
Model Strategy Trainable GSM8K Math
Llama2-7B Bone 87.0M 48.16 8.58
Bat 87.0M 49.36 8.88
RWKV6-7B Bone 55.1M 41.73 6.52
Bat 55.1M 42.76 6.60
In this section, we have verified the effectiveness of non-linear updates by comparing the performance of Bone and Bat, we fine-tuned LLaMA2-7B and RWKV6-7B using both Bone and Bat on the MetaMathQA dataset and evaluated their performance on Math and GSM8K. As shown in the table, with the same number of trainable parameters, the metrics of non-linear updates are better, successfully breaking the original limits.
4.5 Ablation experiments on different grouping methods for Bat
In this subsection, we explore the impact of different grouping methods in the Bat structure on model fine-tuning performance. Due to structural differences in the weight matrix W, the Bat-free grouping requires manual configuration, which is inconvenient. Therefore, this subsection only compares row-wise and column-wise grouping, both of which can be easily extended to any structure. We fine-tuned LLaMA2-7B on the MetaMathQA dataset and validated the results on GSM8k and Math. The results are shown in Table  5. Since in LLaMA2-7B, the dimension of gate_proj in the MLP part is $(4096,11008)$, this leads to an asymmetry between row-wise and column-wise grouping in the Bat structure, making it difficult to align parameter counts. Although Bat-row uses 15M fewer parameters than Bat-col, it still delivers excellent performance. However, this discrepancy in parameter counts makes it challenging to accurately evaluate the differences between the grouping methods.
To explore the differences between the two grouping methods and the effect of parameter count, we added a comparative experiment with RWKV6-3B, as RWKV6’s symmetrical structure ensures that the trainable parameter count is the same whether using row-wise or column-wise grouping during Bat fine-tuning. This allows for a fairer comparison between Bat-row and Bat-col.
Table 5: Comparing Bat-row, Bat-col on math tasks
Model Strategy Trainable GSM8K Math
Llama2-7B Bat-row 72.8M 45.76 7.82
Bat-col 87.0M 49.36 8.88
RWKV6-3B Bat-row 55.1M 25.93 3.12
Bat-col 55.1M 25.25 3.09
The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well. Therefore, we believe that Bat can effectively fit data regardless of the grouping method used. The key factor influencing Bat’s performance remains the block size. However, this doesn’t imply that different grouping methods are meaningless. As more LLMs begin to use techniques like GQA and MLA Liu et al. (2024a) to reduce KV cache overhead Dai et al. (2024); Lee et al. (2024); Shazeer (2019), the main weight matrices become smaller, and Bat will need to adjust its grouping or employ other techniques to adapt to these new technologies.
5 Resource and efficiency
Table  6 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA, Bone, and Bat on a single 4090 GPU. The specific fine-tuning settings are as follows: batch size = 1, context length (ctx_len) = 512.
The results show that Bone has the highest computational efficiency, being nearly 10% faster than LoRA while also being more memory-efficient. However, Bat incurs significantly higher memory usage due to large intermediate values and is slower in comparison.
At the end of the table, we provide the actual resource costs for fine-tuning RWKV6 on the MetaMathQA dataset using 4 NVIDIA 4090 GPUs, with checkpoint techniques applied.
Table 6: Resource and efficiency
Model Strategy Trainable GPU Memory Token throughput
RWKV-3B LoRA 55.1M 12074 MB 3.62 kt/s
Bone 55.1M $\mathbf{11052}$ MB $\mathbf{3.99}$ kt/s
Bat 55.1M 22978 MB 2.16 kt/s
RWKV-3B LoRA(use checkpoint) 55.1M 4*15328 MB 15.6 kt/s
Bone(use checkpoint) 55.1M $\mathbf{4*15304}$ MB $\mathbf{16.0}$ kt/s
Bat(use checkpoint) 55.1M 4*15305 MB 14.2 kt/s
6 Conclusion
In this work, we proposed the DiSHA framework, which divides pre-trained weights into multiple shards and updates them using a shared trainable matrix. This approach significantly reduces resource overhead and opens up new directions for PEFT techniques. Building on the DiSHA framework, we introduced Bone, a broad and efficient method comparable to LoRA. Extensive experiments demonstrated the superiority of Bone, which outperforms LoRA and its variants in evaluation metrics, computational efficiency, and resource usage.
To address the potential issue of linear correlation in shard updates within Bone and the DiSHA framework, we proposed Bat. By applying a linear transformation between the trainable matrix and the pre-trained weights, Bat ensures nonlinear independence in the updates of different shards, successfully breaking through the structural limitations of Bone
7 Future work
Due to the limitations of the experimental equipment, we did not report any results from the full-scale training required for this work. However, through extensive evaluations and comparisons with LoRA and its variants, we validated the effectiveness of DiSHA, Bone, and Bat. The outstanding performance demonstrated in LLMs by this work gives us great confidence. We believe that DiSHA, Bone, and Bat can be applied to any multimodal task. Of course, this work also has some limitations that need to be addressed. Although Bat successfully overcomes the limitations of Bone, it also leads to reduced computational efficiency and increased resource consumption. Therefore, in the future, we plan to develop specialized optimized operators to address this issue.
We welcome the community to provide additional suggestions and conduct further tests.
8 Acknowledgments
I would like to express my heartfelt gratitude to all the seniors who provided valuable feedback on my work. Their insightful suggestions have greatly enhanced the quality of this paper. I also wish to thank RWKVOS for their support. Finally, I sincerely appreciate the meticulous guidance provided by the anonymous online mentor, Professor Cormen, whose help has been indispensable.